40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 10 01:25:28 2020, gautam, Update, LSC, Locking notes 
    Reply  Fri Jul 10 15:26:28 2020, gautam, Update, LSC, MC2 coils need DC balancing? MC2_balancing.pngASDC.png
       Reply  Sat Jul 11 18:24:30 2020, Koji, Update, LSC, MC2 coils need DC balancing? 
       Reply  Sun Jul 12 02:42:01 2020, gautam, Update, LSC, Locking (on rossa) ASCsteps.png
       Reply  Mon Jul 13 11:36:08 2020, rana, Update, LSC, MC2 coils need DC balancing? 
Message ID: 15468     Entry time: Fri Jul 10 15:26:28 2020     In reply to: 15466     Reply to this: 15470   15471   15474
Author: gautam 
Type: Update 
Category: LSC 
Subject: MC2 coils need DC balancing? 

I was looking at some signals from last night, see Attachment #1.

  • It looks like as the DC control signal to the MC2 suspension increases, the MC transmission decreases.
    • I confirmed that the IMC REFL level doesn't correspondingly trend up, but didn't include it here for plot compactness, so I think the cavity length isn't being detuned.
    • So the problem is suggestive of some L2A coupling, and the MC2 coil actuators need to be balanced better at DC?
    • You can see from the IMC WFS control signals that the WFS servo is presumably trying to counteract this L2A action, but doesn't succeed, probably because the servo isn't tuned correctly.
    • This is a problem that is distinct from the drifting TT alignment. So it complicates the alignment situation.
    • Ideally, if the dither alignment servos could be made to work for the arm cavities when locked in the PRFPMI config, this wouldn't be so much of a problem, as the TTs would just adjust the beam pointing to match the cavity axes of the arms. But since I haven't managed to get that servo working yet...
  • But why should MC2 need such a large DC control signal ever?
    • In the PRFPMI lock, the CARM servo is supposed to match the laser frequency to the average length of the two arm cavities.
    • The MC2 suspension is used as a frequency actuator in order to realize this matching.
    • But, as you can see, the digital CARM control signal picks up a significant DC offset the deeper we go into the lock.
    • Can't we offload this DC signal to the laser crystal temperature servo? Is there going to be some weird interaction with the existing slow loop? Or is the idea itself flawed?

Attachment #2 shows some ASC metrics. My conclusion here is that running the PRCL and MICH dither alignment servos (former demodulating REFLDC and latter demodulating ASDC to get an error signal) that running the dither alignment servo and hand tuning the arm ASC loop offsets improves the mode matching to the IFO, because:

  1. The arm transmission increases.
  2. POPDC increases.
  3. ASDC decreases.

The REFLDC behavior needs a bit more interpretation I think, because if the IFO is overcoupled (as I claim it is), then better alignment would at some point actually result in REFLDC increasing. 

All the DC signals recorded by the fast system come from the backplane P2 connector of the PD interface boards. According to the schematic, these signals have a voltage gain of 2. The LSC photodiodes themselves have a nominal DC gain of 50 ohms. So, the conversion from power to digital counts is: 0.8 A/W * 50 V/A * 2 * 3276.8 cts/V * whtGain. Inverting, I get 3.8 uW/ct for a whitening gain of 1. This is power measured at the photodiode - optical losses upstream of the photodiode will have to be accounted for separately.

Assuming a modulation depth of 0.2, the 55 MHz sideband power should be ~20 mW. The Schnupp asymmetry is supposed to give us O(1) transmission of this field to the AS port. Then, the SRM will attenuate the field by a factor of 10, so we expect ~2 mW at the AS port. Let's assume 80 % throughput of this field to the AP table, and then there is a 50/50 beamsplitter dividing the light between the AS55 and AS110 photodiodes. So, we expect there to be ~700 uW of power in the TEM00 mode 55 MHz sideband field. This corresponds to 1600 cts according to the above calibration (the ASDC whitening gain is set to 18 dB). The fact that much smaller numbers were seen for ASDC indicates that (i) the schnupp asymmetry is not so perfectly tuned and the actual transmission of the sideband field to the dark port is smaller, or (ii) one or more optical splitting fractions assumed above is wrong. If the former is true, we can still probably infer the contrast defect if we can somehow get an accurate measurement of the sideband transmission to the dark port.

Attachment 1: MC2_balancing.png  1.934 MB  Uploaded Fri Jul 10 16:27:02 2020  | Hide | Hide all
MC2_balancing.png
Attachment 2: ASDC.png  463 kB  Uploaded Fri Jul 10 17:35:40 2020  | Hide | Hide all
ASDC.png
ELOG V3.1.3-